6. Solve the system r = Ac for

Can you solve this question?
Answer #1

n= An for A = [ 3] characteristic equation of his 1 * =0 = (1-4) (3-4) 41=0 7 m2-44 +420 - 2-2 j2 20 - 1=2,2. Eigen values oNow, AZ 1207 1-11 lo 2 TL-I 1 Nt Son in xet) = g At no est ent no = e .e no 7-1 7- La 7+2 alt) - 1722